OpenStudy (jaedyn):

Diana made a recipe that yields 6 and 1 over 2 cups. If each serving is 1 over 4 cup, which expression will help Diana determine the number of servings her recipe will yield? 6 and 1 over 2 ⋅ 1 over 4 6 and 1 over 2 ÷ 1 over 4 6 and 1 over 2 + 1 over 4 6 and 1 over 2 – 1 over 4

OpenStudy (calculusxy):

1 serving = 1/4 cup We can make an equation to find the relation between the number of servings (s) and the number of cups it produces (c). \(\large \frac{1}{4}s = c\) We want to find out how many servings is produced by \(6 \frac{1}{2}\) cups. \(\large \frac{1}{4}s = 6 \frac{1}{2}\) To find out what \(s\) equals to in that equation, we need to divide both sides by \(\frac{1}{4}\). \[\large \frac{\frac{1}{4}s}{\frac{1}{4}} = \frac{6 \frac{1}{2}}{\frac{1}{4}}\] Therefore \(s\) would equal to \(\large \frac{6 \frac{1}{2}}{\frac{1}{4}}\)

OpenStudy (calculusxy):

@Jaedyn Do you understand?

OpenStudy (jaedyn):

Yes